# StarDist: A Code Generator for Distributed Graph Algorithms

### v1

@@ I Introduction
For distributed graph algorithms to perform reasonably, the cost of communication must compensate that of page faults, critical section overheads and inefficient RAM-processor affinity (NUMA) utilization frequently noted in sequential alternatives.. On compute clusters with SAN storage and RDMA enabled network interfaces, network communication is often faster than storage[8].
IO intensive workloads on large data, give distributed algorithms a scope to outperform shared memory or sequential algorithms. The time, otherwise spent on page faults, would be utilized in computing a partial-result on another node, and the communication of the computed result would be faster than a fault. Graph algorithms are IO intensive, due to their irregular memory access patterns. Even for reasonably sized graphs, the cache miss rate is considerable[9].
@@
    A bulk-reduction substrate backed by passive MPI RMA calls in junction with a distributed work list. This eliminates the need for receiving ranks to be aware of ranks and sizes of incoming packets and also exploits RDMA capabilities of underlying hardware[2].
@@
    We define reduction-exclusive statements that provide a stub for achieving optimal communication patterns. 1b is an example where both iterations are reduction-exclusive. Reduction-exclusive statements allow delayed updates enabling our work to perform on par with Gluon[1] used in Galois [3] which utilize synchronization passes for synchronizing master and mirror nodes.
@@ II Related Works
Distributed Galois[1] provides a framework leveraging gluon-async[1] that can be used to write peer-to-peer distributed graph algorithms. Gluon Async[1] provides a bulk asynchronous parallel model which provides mechanisms to bypass global locks on distributed memory and allows passive updates across participants through proxy nodes.
@@
Gemini [4] framework adapts to runtime and provides an effective edge centric graph partitioning scheme. It facilitates writing graph algorithms both in shared and distributed memory situations. Much like gluon async [1] in distributed Galois [3], updates are propagated iteratively until convergence.
DRONE[2] explores a subgraph boundary synchronization scheme over a vertex-cut based partitioning, designed similar to d-Galois but with the additional clause of only boundary nodes having master and mirror nodes and a master-slave architecture where the master node responsible for synchronization and book-keeping. DRONE [2] claims better performance over gluon[1] and gemini [4]
In all three of the above frameworks [1] [4] [2], the propagation of updates to mirror nodes and their synchronization happens in pulses, as defined in Goldberg and Tarjan’s parallel implementation of push relabel[8]. The convergence paradigm represents an inherent universality as it works across shared memory multi-core, distributed memory, as well as GPUs. However, for distributed Graph algorithms, every pulse is expensive and bulkier pulses happening less frequently achieving similar results is a faster directive.
Poly[10] The exploitation of reduction operations in polyhedral analysis are detailed in [10] where multiple possible reduction operations were evaluated to perform 2.21 times faster than control. Semantic knowledge of delay-able statements present in reduction operators allowed scheduling of multi-level iterations without both memory and control barrier delays.
Cagra[9] The irregular memory access patterns with graph algorithms and their effects when scaled are studied in Cagra[9], that attempts to bring down the cache miss rate by vertex relabeling. Cagra [9] does a best effort service to fit the neighborhood of a vertex into the L1/L2 cache of the CPU for shared memory frameworks. Distributed memory, especially MPI, would have one process affined to a core, and the vertices managed by a core would be a dividend of the total vertices present in the graph, thus achieving a similar effect with cache misses, in a natural fashion.
StarPlat[5] a versatile graph DSL provides a distributed backend for generating graph algorithms in openMPI, a more detailed explanation of it is within section III. Whereas inclusion of RMA to StarPlat in [6] gave flexibility to code generation, the performance regressed multiple times, small and frequent RMA updates are expensive due to RMA read/update cycle initiation and termination overheads.
The StarPlat compiler’s[6] distributed backend utilizes a peer-to-peer graph data structure in its header library and generates the openMPI code from the DSL code using StarPlat’s graph AST. In this section we describe essential structures in the StarPlat’s distributed backend.
@@
  • Graph Data Structure: The graph partitions, as shown in Figure 2a, in Starplat[5] do not contain ghost or mirror nodes used in gluon-async [1] or DRONE [2]. Updates are directly and instantaneously propagated, and the pulses used to propagate updates in [1], [2] and [4] are not implemented so that the programmer using StarPlat does writes their code without having any awareness of parallel programming. This results in frequent updates that are smaller in size (usually 4 Byte integers).
    The overhead of initiating and terminating RMA cycles for frequent, small transmissions, is not handled by openMPI[7]. Therefore, the performance dipped for static Single Source Shortest Paths from [5] to [6] from 3411s to 6331.941s, an increase of 1.856 times.
@@ III Motivation
The StarPlat compiler’s distributed backend prefers generality in code generation over performance. By profiling generated code for Single Source Shortest Paths using gprof [11], perf and hpctoolkit[12], we deduced function calls that were utilizing high number of CPU cycles. We logged times of these functions and concluded the following phases to be expensive:
@@
The definition of a pulse 3 is an extended, more abstract form of what appears in [8], where delayed and aggregated communication was found beneficial for distributed algorithms.
@@ IV The Backend Analyzer
To adhere to the peer-to-peer and no mirror node ideology, StarPlat initiates communication with owner nodes and transfers the neighborhood information on demand. We noticed through gprof[11], that get_edge calls within neighborhood traversals were taking up a high number of CPU cycles.
@@
OpenMPI[7] does not internally optimize subsequent RMA calls by keeping RMA phases active. Keeping the RMA Window under a shared lock also does not optimize the get calls. We use lemma 1 and our analysis for determining reduction-exclusive statements as in definition 1 to argue that get calls within statements deemed reduction-exclusive can safely de-reference the local memory of an RMA window, as updates to the data undergoing reduction are not going to be propagated until the end of the pulse. We show our analysis in Figure 6

### v2

@@ I Introduction
For distributed graph algorithms to perform reasonably, the cost of communication must compensate that of page faults, critical section overheads and inefficient RAM-processor affinity (NUMA) utilization frequently noted in sequential alternatives.. On compute clusters with SAN storage and RDMA enabled network interfaces, network communication is often faster than storage.
IO intensive workloads on large data, give distributed algorithms a scope to outperform shared memory or sequential algorithms. The time, otherwise spent on page faults, would be utilized in computing a partial-result on another node, and the communication of the computed result would be faster than a fault. Graph algorithms are IO intensive, due to their irregular memory access patterns. Even for reasonably sized graphs, the cache miss rate is considerable[ref9].
@@
    A bulk-reduction substrate backed by passive MPI RMA calls in junction with a distributed work list. This eliminates the need for receiving ranks to be aware of ranks and sizes of incoming packets and also exploits RDMA capabilities of underlying hardware[ref2].
@@
    We define reduction-exclusive statements that provide a stub for achieving optimal communication patterns. 1b is an example where both iterations are reduction-exclusive. Reduction-exclusive statements allow delayed updates enabling our work to perform on par with Gluon[ref1] used in Galois [ref3] which utilize synchronization passes for synchronizing master and mirror nodes.
@@ II Related Works
Distributed Galois[ref1] provides a framework leveraging gluon-async[ref1] that can be used to write peer-to-peer distributed graph algorithms. Gluon Async[ref1] provides a bulk asynchronous parallel model which provides mechanisms to bypass global locks on distributed memory and allows passive updates across participants through proxy nodes.
@@
Gemini [ref4] framework adapts to runtime and provides an effective edge centric graph partitioning scheme. It facilitates writing graph algorithms both in shared and distributed memory situations. Much like gluon async [ref1] in distributed Galois [ref3], updates are propagated iteratively until convergence.
DRONE[ref2] explores a subgraph boundary synchronization scheme over a vertex-cut based partitioning, designed similar to d-Galois but with the additional clause of only boundary nodes having master and mirror nodes and a master-slave architecture where the master node responsible for synchronization and book-keeping. DRONE [ref2] claims better performance over gluon[ref1] and gemini [ref4]
In all three of the above frameworks [ref1] [ref4] [ref2], the propagation of updates to mirror nodes and their synchronization happens in pulses, as defined in Goldberg and Tarjan’s parallel implementation of push relabel[ref8]. The convergence paradigm represents an inherent universality as it works across shared memory multi-core, distributed memory, as well as GPUs. However, for distributed Graph algorithms, every pulse is expensive and bulkier pulses happening less frequently achieving similar results is a faster directive.
Poly[ref10] The exploitation of reduction operations in polyhedral analysis are detailed in [ref10] where multiple possible reduction operations were evaluated to perform 2.21 times faster than control. Semantic knowledge of delay-able statements present in reduction operators allowed scheduling of multi-level iterations without both memory and control barrier delays.
Cagra[ref9] The irregular memory access patterns with graph algorithms and their effects when scaled are studied in Cagra[ref9], that attempts to bring down the cache miss rate by vertex relabeling. Cagra [ref9] does a best effort service to fit the neighborhood of a vertex into the L1/L2 cache of the CPU for shared memory frameworks. Distributed memory, especially MPI, would have one process affined to a core, and the vertices managed by a core would be a dividend of the total vertices present in the graph, thus achieving a similar effect with cache misses, in a natural fashion.
StarPlat[ref5] a versatile graph DSL provides a distributed backend for generating graph algorithms in openMPI, a more detailed explanation of it is within section III. Whereas inclusion of RMA to StarPlat in [ref6] gave flexibility to code generation, the performance regressed multiple times, small and frequent RMA updates are expensive due to RMA read/update cycle initiation and termination overheads.
The StarPlat compiler’s[ref6] distributed backend utilizes a peer-to-peer graph data structure in its header library and generates the openMPI code from the DSL code using StarPlat’s graph AST. In this section we describe essential structures in the StarPlat’s distributed backend.
@@
  • Graph Data Structure: The graph partitions, as shown in Figure 2a, in Starplat[ref5] do not contain ghost or mirror nodes used in gluon-async [ref1] or DRONE [ref2]. Updates are directly and instantaneously propagated, and the pulses used to propagate updates in [ref1], [ref2] and [ref4] are not implemented so that the programmer using StarPlat does writes their code without having any awareness of parallel programming. This results in frequent updates that are smaller in size (usually 4 Byte integers).
    The overhead of initiating and terminating RMA cycles for frequent, small transmissions, is not handled by openMPI[ref7]. Therefore, the performance dipped for static Single Source Shortest Paths from [ref5] to [ref6] from 3411s to 6331.941s, an increase of 1.856 times.
@@ III Motivation
The StarPlat compiler’s distributed backend prefers generality in code generation over performance. By profiling generated code for Single Source Shortest Paths using gprof [ref11], perf and hpctoolkit[ref12], we deduced function calls that were utilizing high number of CPU cycles. We logged times of these functions and concluded the following phases to be expensive:
@@
The definition of a pulse 3 is an extended, more abstract form of what appears in [ref8], where delayed and aggregated communication was found beneficial for distributed algorithms.
@@ IV The Backend Analyzer
To adhere to the peer-to-peer and no mirror node ideology, StarPlat initiates communication with owner nodes and transfers the neighborhood information on demand. We noticed through gprof[ref11], that get_edge calls within neighborhood traversals were taking up a high number of CPU cycles.
@@
OpenMPI[ref7] does not internally optimize subsequent RMA calls by keeping RMA phases active. Keeping the RMA Window under a shared lock also does not optimize the get calls. We use lemma 1 and our analysis for determining reduction-exclusive statements as in definition 1 to argue that get calls within statements deemed reduction-exclusive can safely de-reference the local memory of an RMA window, as updates to the data undergoing reduction are not going to be propagated until the end of the pulse. We show our analysis in Figure 6